# Quasinormal modes of regular black holes surrounded by skewed dark matter distributions

### v1

@@ 1 Introduction
Regular black holes have emerged as one of the most active research topics in the contemporary black hole physics [22, 23, 24, 25, 26, 27]. The term ‘regular’ refers to black holes that are free of central curvature singularities. The absence of a singularity is typically attributed to the presence of a core at the center of regular black holes that violates the strong energy condition, thereby evading the Penrose singularity theorem and effectively removing the singularity [28, 29].
In this work, we focus exclusively on a class of recently proposed regular black holes that arises from the gravitational collapse of matter [22, 30, 31, 32], supplemented by quantum corrections, such as those predicted by asymptotic safety [22] and loop quantum gravity theories [30]. These models are of particular interest not only because they provide a concrete mechanism for the formation of regular black holes, but more importantly, because their construction implies that the spacetime outside horizons can carry imprints of their interior structure. In this class of regular black holes, the interior spacetime is first determined by the collapse of matter, including possible quantum corrections, and the exterior metric is then derived via appropriate matching conditions [33, 34, 35]. As a result, such models offer a potential window into probing the otherwise inaccessible spacetime inside event horizons.
Moreover, we employ the quasinormal mode (QNM) analysis [36] to study the impact of dark matter on the gravitational waves emitted by regular black holes. QNMs refer to damped oscillation modes that characterize the waveform of ringdown phases following binary black hole mergers [37]. The QNM analysis is widely used in the current study of regular black holes[38, 39, 40, 41, 42], most of which focus, however, on vacuum solutions, leaving matter effects on singular black holes. The recent research shows that the presence of matter modifies the effective potential in the perturbation equations [43, 44, 45, 46, 47], introducing a second potential barrier or well, which in turn affects the stability of perturbations. Our work extends the research by incorporating a skewed dark matter distribution around regular black holes, revealing novel features, such as long-lived modes and enhanced polar perturbation sensitivity, which would probably inform future gravitational wave observations.
@@
The structure of this paper is as follows. In Sec. 2, we begin by analyzing the ringdown waveform of the model without dark matter [22]. When the parameter $\xi$ approaches zero, the model reduces exactly to the Schwarzschild black hole, meaning that $\xi$ characterizes the deviation from Schwarzschild geometry. Our analysis reveals that the ringdown waveform exhibits distinct features from those of the Schwarzschild black hole, particularly during the initial phase. [^1] Our classification of waveforms follows the conventions established in Refs. [37, 48] (see Fig. 3 in either reference). The waveform is typically divided into three distinct stages: The initial transient phase, the ringdown phase, and the late-time tail phase.
@@ 2 Quasinormal modes without dark matter
In this section, we revisit the QNMs of a regular black hole derived from the framework of asymptotically safe gravity. Unlike Ref. [49], which focused on the QNM spectrum and gray-body factors, we make our analysis from the perspective of waveform evolution. This distinction ensures that our investigation is not a repetition of previous calculations. Our primary goal here is to establish a complete baseline for comparison, particularly with the dark-matter-modified model introduced in the following sections.
The black hole model under consideration is formed via dust collapse, with quantum corrections inspired by asymptotic safety. Its exterior spacetime metric takes [22] the form,
@@
while the radius of the ISCO, defined by the conditions $V_{\rm eff}^{\prime}=0=V_{\rm eff}^{\prime\prime}$, gives $r_{\rm ISCO}\approx 5.51M$. Here, $V_{\rm eff}=f(r)\left(\frac{L^{2}}{r^{2}}+1\right)$ refers to the effective potential for a massive test particle moving along geodesics [50, 51], where $L$ denotes the angular momentum of the test particle. To distinguish it from the effective potential in the perturbation equations discussed later, we explicitly denote the former with the subscript “eff” and use the calligraphic symbol $\mathcal{V}$ for the latter without subscripts. We mention in advance that the location of the ISCO will play a key role in our modeling of the dark matter distribution in the subsequent sections.
We now examine the perturbations of three types of test fields in the single-horizon background. The corresponding effective potentials of the master equations take [49] the forms,
@@
$\mathcal{V}^{(d,\pm)}(r)=f(r)\frac{\ell+1}{r}\left[\frac{\ell+1}{r}\mp\sqrt{f(r)}\mp\frac{\mathrm{d}}{\mathrm{d}r}\sqrt{f(r)}\right],$ (5c)
[…]
@@
[…]
The ringdown waveform is shown in Fig. 2, where the initial perturbation is modeled as a Gaussian wave packet. [^3] The ringdown waveform is computed using a null coordinate $(u,v)$ grid discretization of the wave equation with a four-point finite difference stencil, achieving $O(h^{4})$ local accuracy and systematic integration along null characteristics [54, 52]. Due to the structural similarity of the effective potentials, panels 2(a), 2(b), and 2(d) exhibit similar waveform features: The signal initially displays damped oscillations, and then gradually diminishes into a monotonic decay. As the angular quantum number $\ell$ increases, both the oscillation frequency and the decay rate increase, corresponding to a larger real and minus imaginary parts of QNM frequencies, respectively, which results in a longer-lasting oscillation and a faster subsequent damping.
@@
Panel 2(c), as predicted from the structure of the effective potential, exhibits a qualitatively different behavior: The late-time tail waveform grows exponentially over time, signaling an instability. This divergence is consistent with the presence of a potential well beyond the main barrier. Moreover, as $\ell$ increases from 111 to 222 and to 333, the absolute value of barrier peak-to-well depth ratios grows monotonically from 3.413.413.41 to 6.286.286.28 and to 10.5410.5410.54. This indicates that the influence of the potential well weakens progressively, and accordingly, the onset of the late-time tail divergence is delayed.
@@ 3.1 Introducing dark matter
where $\rho_{0}$ is a parameter with density dimension, $r_{0}\gtrsim r_{\rm ISCO}$, [^4] Actually, the presence of a radial-dependent mass correction inevitably shifts the location of the ISCO in the modified model relative to that in the unmodified case. $\sigma$ is the width, and $\alpha>0$ is shape parameter. In our analysis, the peak position $r_{0}$ of the dark matter distribution is chosen to lie outside the ISCO, motivated by considerations of environmental stability. If dark matter were distributed inside the ISCO, it would continuously fall into the merged black hole, leading to a time-dependent evolution of the spacetime and corresponding modifications to the effective potential. Such dynamical scenarios are inherently more complex and are not addressed in the present work. In contrast, when $r_{0}>r_{\rm ISCO}$, the dark matter configuration can be reasonably treated as quasi-stationary, resulting in a static modification to the effective potential.
The skewed normal distribution implies that the dark matter density is high near black holes, but low away from black holes. This assumption comes from the strong gravity of black holes, which is fully consistent with physically realistic scenarios. It is worth emphasizing that there exist many possible implementations of skewed normal distributions. In this work, we adopt a representative form, namely, a modified Gaussian profile with skewness, to illustrate the essential physical features without loss of generality. In addition, the mass density profile introduced in Eq. (6) provides an opportunity for comparison with the well-established Einasto profile [55, 56], which is commonly used in modeling dark matter halos,
@@
as given by Eq. (6). It should be emphasized that our analysis addresses only the clustering of dark matter around regular black holes, consistent with recent findings based on binary companion orbital decay, such as Ref. [57, 58]. We do not consider scenarios in which dark matter itself emits gravitational waves spontaneously (e.g., via ultralight boson annihilation forming scalar clouds), as examined in e.g. the LIGO–Virgo–KAGRA Collaboration’s search for long lasting and quasi-monochromatic signals[59]. Rather, our focus is strictly on how the ringdown-phase gravitational wave signature is altered when a regular black hole is embedded in a dark matter environment.
@@ 3.2 Quasinormal mode analysis
Thus, the primary effect of introducing a skewed normal correction to the mass profile is the formation of a relatively small potential well in the Type I effective potential, and it gives rise to a narrower well in the Type II effective potential. This behavior is qualitatively different from that studied in Ref. [44, 46], where a direct Gaussian modification to the effective potential produced a barrier-like correction instead.
@@ 4 Differential sensitivity of axial and polar tensor field perturbations
[…]
@@
We now incorporate dark matter into our analysis. Following the approach in Ref. [21], we neglect the backreaction of the energy-momentum tensor in the perturbation equations. This approximation simplifies the mathematical treatment but remains valid under the assumption of a small density parameter $\rho_{0}$, which ensures that higher-order corrections from the stress-energy tensor are suppressed. Although this neglect may introduce systematic errors in the effective potential, particularly for configurations with dense dark matter distributions, the resulting framework provides a tractable description of the system’s dynamics.
We consider a skewed normal distribution for the dark matter density profile, as specified in Eq. (6). The corresponding effective potentials are presented in Figs. 6(a) and 6(b) for axial and polar perturbations, respectively. As we shall demonstrate below, the presence of dark matter breaks the symmetry between axial and polar modes that characterizes the vacuum case, leading to qualitatively different perturbation dynamics. A crucial observation emerges from our analysis: The introduction of dark matter produces barrier modifications rather than potential well modifications, contrasting sharply with the test field scenarios examined in the previous section. This modification to the main potential barrier is expected to induce echo-like features in the resulting waveform. Such a behavior can be anticipated from the emergence of a double-peak structure in the effective potential. These configurations act effectively as cavities, partially reflect and trap gravitational wave energy between successive potential barriers, and ultimately generate late-time echo signatures. This mechanism aligns with the previous analyses of gravitational wave echoes in modified gravity scenarios [45].
@@ 5 Discussion and outlook
Finally, as gravitational wave detectors improve in precision, it becomes increasingly tangible for us to detect deviations from singular black hole signatures due to surrounding dark matter distributions or the absence of a central singularity. Our study contributes to this growing effort by providing a theoretically consistent framework for modeling such deviations in the context of regular black holes surrounded by dark matter. One of urgent issues that needs to be solved is to advance this framework toward more realistic astrophysical scenarios, i.e., to extend our analyses to rotating regular black holes. In such cases, the introduction of dark matter, especially with a skewed distribution, could interact nontrivially with the frame-dragging effects of rotation, potentially leading to qualitatively different modifications of ringdown waveforms. To model such systems, one promising idea is to employ the extended Newman–Janis algorithm [61] or its generalizations to construct rotating counterparts of regular black holes with dark matter corrections. It will be a central focus of our future research to investigate the ringdown behavior in these rotating spacetimes and analyze the axial–polar mode splitting, damping times, and possible echoes.
@@ Acknowledgements
C.L. and Y.-L. T. have an equal contribution. This work was supported in part by the National Natural Science Foundation of China under Grant No. 12175108. Moreover, L.C. is supported by Yantai University under Grant No. WL22B224 and Z.-X. Z is supported by the Pilot Scheme of Talent Training in Basic Sciences (Boling Class of Physics, Nankai University), Ministry of Education.
@@ Appendix A Schwarzschild black hole surrounded by dark matter
The emergence of double-barrier structures in the effective potentials provides the necessary conditions for gravitational wave echo generation [62]. However, the observation of distinct echo signals requires specific geometric constraints on the potential profile, particularly regarding the separation distance between successive barriers. In our configuration, while the barrier structures are present, the relatively modest separation distances suggest that clear echo identification would require highly sensitive gravitational wave detectors and sophisticated data analysis techniques.

### v2

@@ 1 Introduction
Regular black holes have emerged as one of the most active research topics in the contemporary black hole physics [Bonanno:2023rzk, Bueno:2024dgm, Bueno:2024eig, Bronnikov:2024izh, Bueno:2024zsx, Carballo-Rubio:2022nuj]. The term ‘regular’ refers to black holes that are free of spacetime singularities, i.e., that are typically characterized by geodesic completeness as per the singularity theorems. Moreover, this term is often accompanied by curvature finiteness in spacetime. [^1] However, these two definitions are not always consistent. For a detailed discussion of this issue, we refer the reader to Ref. [Zhang:2025nim]. The absence of a singularity is typically attributed to the presence of a core at the center of regular black holes that violates the strong energy condition, thereby evading the Penrose singularity theorem and effectively removing the singularity [Bambi:2023try, Lan:2023cvz].
In this work, we focus exclusively on a class of recently proposed regular black holes that arises from the gravitational collapse of matter [Bonanno:2023rzk, Cipriani:2024nhx, Vertogradov:2025yto, Vertogradov:2025snh], supplemented by quantum corrections, such as those predicted by asymptotic safety [Bonanno:2023rzk] and loop quantum gravity theories [Cipriani:2024nhx]. These models are of particular interest not only because they provide a concrete mechanism for the formation of regular black holes, but more importantly, because their construction implies that the spacetime outside horizons can carry imprints of their interior structure. In this class of regular black holes, the interior spacetime is first determined by the collapse of matter, including possible quantum corrections, and the exterior metric is then derived via appropriate matching conditions [Israel:1966rt, Barrabes:1991ng, Fayos:1996gw]. As a result, such models offer a potential window into probing the otherwise inaccessible spacetime inside event horizons.
Moreover, we employ the quasinormal mode (QNM) analysis [Berti:2025hly] to study the impact of dark matter on the gravitational waves emitted by regular black holes. QNMs refer to damped oscillation modes that characterize the waveform of ringdown phases following binary black hole mergers [Konoplya:2011qq]. The QNM analysis is widely used in the current study of regular black holes[Konoplya:2022hll, Konoplya:2023aph, Konoplya:2023ahd, Konoplya:2025hgp, Lan:2023vaa], most of which focus, however, on vacuum solutions, leaving matter effects on singular black holes. The recent research shows that the presence of matter modifies the effective potential in the perturbation equations [Li:2019kwa, Cheung:2021bol, Huang:2021qwe, Berti:2022xfj, Cao:2024sot, Mai:2025cva], introducing a second potential barrier or well, which in turn affects the stability of perturbations. Our work extends the research by incorporating a skewed dark matter distribution around regular black holes, revealing novel features, such as long-lived modes and enhanced polar perturbation sensitivity, which would probably inform future gravitational wave observations.
@@
The structure of this paper is as follows. In Sec. 2, we begin by analyzing the ringdown waveform of the model without dark matter [Bonanno:2023rzk]. When the parameter $\xi$ approaches zero, the model reduces exactly to the Schwarzschild black hole, meaning that $\xi$ characterizes the deviation from Schwarzschild geometry. Our analysis reveals that the ringdown waveform exhibits distinct features from those of the Schwarzschild black hole, particularly during the initial phase. [^2] Our classification of waveforms follows the conventions established in Refs. [Konoplya:2011qq, Burgess:2023pny] (see Fig. 3 in either reference). The waveform is typically divided into three distinct stages: The initial transient phase, the ringdown phase, and the late-time tail phase.
@@ 2 Quasinormal modes without dark matter
In this section, we revisit the QNMs of a regular black hole derived from the framework of asymptotically safe gravity. Unlike Ref. [Stashko:2024wuq], which focused on the QNM spectrum and gray-body factors, we make our analysis from the perspective of waveform evolution. This distinction ensures that our investigation is not a repetition of previous calculations. Our primary goal here is to establish a complete baseline for comparison, particularly with the dark-matter-modified model introduced in the following sections.
The black hole model under consideration is formed via dust collapse, with quantum corrections inspired by asymptotic safety. Its exterior spacetime metric takes [Bonanno:2023rzk] the form,
@@
while the radius of the ISCO, defined by the conditions $V_{\rm eff}^{\prime}=0=V_{\rm eff}^{\prime\prime}$, gives $r_{\rm ISCO}\approx 5.51M$. Here, $V_{\rm eff}=f(r)\left(\frac{L^{2}}{r^{2}}+1\right)$ refers to the effective potential for a massive test particle moving along geodesics [Chandrasekhar:1985kt, Guo:2025scs], where $L$ denotes the angular momentum of the test particle. To distinguish it from the effective potential in the perturbation equations discussed later, we explicitly denote the former with the subscript “eff” and use the calligraphic symbol $\mathcal{V}$ for the latter without subscripts. We mention in advance that the location of the ISCO will play a key role in our modeling of the dark matter distribution in the subsequent sections.
We now examine the perturbations of three types of test fields in the single-horizon background. The perturbations of massless test fields in this background satisfy master equations of the form $\frac{d^{2}\Psi}{dr_{*}^{2}}+(\omega^{2}-V(r))\Psi=0$, where $r_{*}$ is the tortoise coordinate, $\omega$ is the mode frequency, and $\Psi$ is the radial wave function [Chandrasekhar:1985kt, Pani:2013pma]. These equations are derived from the Klein-Gordon equation for scalar fields, the Maxwell equations for vector fields, and the Dirac equation for spinor fields, after variable separation and appropriate rescaling. The corresponding effective potentials are [Zinhailo:2019rwd, Stashko:2024wuq]
@@
$\mathcal{V}^{(d,\pm)}(r)=f(r)\frac{\ell+1}{r}\left[\frac{\ell+1}{r}\mp\sqrt{f(r)}\pm\frac{\mathrm{d}}{\mathrm{d}r}\sqrt{f(r)}\right],$ (5c)
where the superscripts $(s)$, $(v)$, and $(d,\pm)$ on the potential $\mathcal{V}(r)$ label scalar, vector, and spinor field perturbations, respectively. The tortoise coordinate is defined via $r^{*}=\int\frac{\mathrm{d}r}{f(r)}$. The corresponding effective potentials [^3] It is important to note that for positive-parity spinor field perturbations, if $\ell=0$, the effective potential takes the form of a pure potential well, whereas if $\ell>1$, it becomes a pure potential barrier. A similar behavior is also observed in the case of axial tensor field perturbations. are illustrated in Fig. 1.
@@
As shown in Fig. 1, the potentials exhibit standard bell-shaped profiles for the scalar $(s)$, vector $(v)$, and negative-parity spinor (d,)(d,-) field perturbations. However, the positive-parity spinor field perturbation $(d,+)$ displays a qualitatively different structure, presenting a potential well located to the right of the main barrier. According to our previous studies [Zhang:2025dzt, Lan:2025cvs], this kind of effective potential structures can give rise to unstable QNMs, meaning that the associated perturbation waveform grows unbounded over time. The onset of such instability is determined by the depth of the potential well, particularly the ratio between the peak of the potential barrier and the depth of the well.
The ringdown waveform is shown in Fig. 2, where the initial perturbation is modeled as a Gaussian wave packet. [^4] The ringdown waveform is computed using a null coordinate $(u,v)$ grid discretization of the wave equation with a four-point finite difference stencil, achieving $O(h^{4})$ local accuracy and systematic integration along null characteristics [Gundlach:1993tp, Zhang:2025dzt]. Due to the structural similarity of the effective potentials, panels 2(a), 2(b), and 2(d) exhibit similar waveform features: The signal initially displays damped oscillations, and then gradually diminishes into a monotonic decay. As the angular quantum number $\ell$ increases, both the oscillation frequency and the decay rate increase, corresponding to a larger real and minus imaginary parts of QNM frequencies, respectively, which results in a longer-lasting oscillation and a faster subsequent damping.
@@
Panel 2(c), as predicted from the structure of the effective potential, exhibits a qualitatively different behavior: The late-time tail waveform grows exponentially over time, signaling an instability. This divergence is consistent with the presence of a potential well beyond the main barrier. Moreover, as $\ell$ increases from $1$ to $2$ and to $3$, the absolute value of barrier peak-to-well depth ratios grows monotonically from $3.41$ to $6.28$ and to $10.54$. This indicates that the influence of the potential well weakens progressively, and accordingly, the onset of the late-time tail divergence is delayed.
@@ 3.1 Introducing dark matter
where $\rho_{0}$ is a parameter with density dimension, $r_{0}\gtrsim r_{\rm ISCO}$, [^5] Actually, the presence of a radial-dependent mass correction inevitably shifts the location of the ISCO in the modified model relative to that in the unmodified case. $\sigma$ is the width, and $\alpha>0$ is shape parameter. This density profile is a phenomenological choice inspired by skewed distributions in accretion disks and dark matter halos, where the asymmetry arises from gravitational clustering [Binney:2008gds, Navarro:2003ew, Springel:2008cc]. It resembles the matter configuration in standard thin-disk models but with skewness to reflect higher densities near the black hole. In our analysis, the peak position $r_{0}$ of the dark matter distribution is chosen to lie outside the ISCO, motivated by considerations of environmental stability. If dark matter were distributed inside the ISCO, it would continuously fall into the merged black hole, leading to a time-dependent evolution of the spacetime and corresponding modifications to the effective potential. Such dynamical scenarios are inherently more complex and are not addressed in the present work. In contrast, when $r_{0}>r_{\rm ISCO}$, the dark matter configuration can be reasonably treated as quasi-stationary, resulting in a static modification to the effective potential.
The skewed normal distribution implies that the dark matter density is high near black holes, but low away from black holes. This assumption comes from the strong gravity of black holes, which is fully consistent with physically realistic scenarios. It is worth emphasizing that there exist many possible implementations of skewed normal distributions. In this work, we adopt a representative form, namely, a modified Gaussian profile with skewness, to illustrate the essential physical features without loss of generality. In addition, the mass density profile introduced in Eq. (6) provides an opportunity for comparison with the well-established Einasto profile [Navarro:2003ew, Graham:2005xx], which is commonly used in modeling dark matter halos,
@@
as given by Eq. (6). It should be emphasized that our analysis addresses only the clustering of dark matter around regular black holes, consistent with recent findings based on binary companion orbital decay, such as Ref. [Chan:2022gqd, Belikov:2013nca]. We do not consider scenarios in which dark matter itself emits gravitational waves spontaneously (e.g., via ultralight boson annihilation forming scalar clouds), as examined in e.g. the LIGO–Virgo–KAGRA Collaboration’s search for long lasting and quasi-monochromatic signals[LIGOScientific:2021rnv]. Rather, our focus is strictly on how the ringdown-phase gravitational wave signature is altered when a regular black hole is embedded in a dark matter environment.
@@ 3.2 Quasinormal mode analysis
Thus, the primary effect of introducing a skewed normal correction to the mass profile is the formation of a relatively small potential well in the Type I effective potential, and it gives rise to a narrower well in the Type II effective potential. This behavior is qualitatively different from that studied in Ref. [Cheung:2021bol, Berti:2022xfj], where a direct Gaussian modification to the effective potential produced a barrier-like correction instead.
@@ 4 Differential sensitivity of axial and polar tensor field perturbations
The effective potentials for axial $(a)$ and polar $(p)$ tensor field perturbations are given by [del-Corral:2022kbk]
@@
We now incorporate dark matter into our analysis. Following the approach in Ref. [Figueiredo:2023gas], we neglect the backreaction of the energy-momentum tensor in the perturbation equations. This approximation simplifies the mathematical treatment but remains valid under the assumption of a small density parameter $\rho_{0}$, which ensures that higher-order corrections from the stress-energy tensor are suppressed. Although this neglect may introduce systematic errors in the effective potential, particularly for configurations with dense dark matter distributions, the resulting framework provides a tractable description of the system’s dynamics.
We consider a skewed normal distribution for the dark matter density profile, as specified in Eq. (6). The corresponding effective potentials are presented in Figs. 6(a) and 6(b) for axial and polar perturbations, respectively. As we shall demonstrate below, the presence of dark matter breaks the symmetry between axial and polar modes that characterizes the vacuum case, leading to qualitatively different perturbation dynamics. A crucial observation emerges from our analysis: The introduction of dark matter produces barrier modifications rather than potential well modifications, contrasting sharply with the test field scenarios examined in the previous section. This modification to the main potential barrier is expected to induce echo-like features in the resulting waveform. Such a behavior can be anticipated from the emergence of a double-peak structure in the effective potential. These configurations act effectively as cavities, partially reflect and trap gravitational wave energy between successive potential barriers, and ultimately generate late-time echo signatures. This mechanism aligns with the previous analyses of gravitational wave echoes in modified gravity scenarios [Huang:2021qwe].
@@ 5 Discussion and outlook
Finally, as gravitational wave detectors improve in precision, it becomes increasingly tangible for us to detect deviations from singular black hole signatures due to surrounding dark matter distributions or the absence of a central singularity. Our study contributes to this growing effort by providing a theoretically consistent framework for modeling such deviations in the context of regular black holes surrounded by dark matter. One of urgent issues that needs to be solved is to advance this framework toward more realistic astrophysical scenarios, i.e., to extend our analyses to rotating regular black holes. In such cases, the introduction of dark matter, especially with a skewed distribution, could interact nontrivially with the frame-dragging effects of rotation, potentially leading to qualitatively different modifications of ringdown waveforms. To model such systems, one promising idea is to employ the extended Newman–Janis algorithm [Lan:2024wfo] or its generalizations to construct rotating counterparts of regular black holes with dark matter corrections. It will be a central focus of our future research to investigate the ringdown behavior in these rotating spacetimes and analyze the axial–polar mode splitting, damping times, and possible echoes.
@@ Acknowledgements
C.L. and Y.-L. T. make an equal contribution. The authors thank Alexander Zhidenko for pointing out a typographical error in the formula. This work was supported in part by the National Natural Science Foundation of China under Grant No. 12175108. Moreover, L.C. is supported by Yantai University under Grant No. WL22B224 and Z.-X. Z is supported by the Pilot Scheme of Talent Training in Basic Sciences (Boling Class of Physics, Nankai University), Ministry of Education.
During the preparation of this work the authors used ChatGPT in order to create the schematic diagram shown in Figure 3. After using this tool/service, the authors reviewed and edited the content as needed and took full responsibility for the content of the published article.
@@ Appendix A Schwarzschild black hole surrounded by dark matter
The emergence of double-barrier structures in the effective potentials provides the necessary conditions for gravitational wave echo generation [Yang:2024prm]. However, the observation of distinct echo signals requires specific geometric constraints on the potential profile, particularly regarding the separation distance between successive barriers. In our configuration, while the barrier structures are present, the relatively modest separation distances suggest that clear echo identification would require highly sensitive gravitational wave detectors and sophisticated data analysis techniques.